1. Find the variation constant and an equation of variation where y varies directly as x and y=12 when x=6.
2. Divide... (35b^3+18b^2+37b+41)/(5b+4)
3. Solve for x... 4x(x-1)-5x(x)=3
4. A 15 in. TV set also has a width of 12 inches. What is it's height?

Answer by stanbon(75887)

1. Find the variation constant and an equation of variation where y varies directly as x and y=12 when x=6.
Direct Variation: y = kx
12 = k*6
k = 2
-------------------------------------------------
2. Divide... (35b^3+18b^2+37b+41)/(5b+4)
Use synthetic division:
-4/5)....35....18....37....41
...........35....-10..45...|..5
Quotient: 35x^2 -10x + 45
Remainder: -5
---------------------------------
3. Solve for x... 4x(x-1)-5x(x)=3
-x^2 -4x -3 = 0
x^2 + 4x + 3 = 0
(x+3)(x+1) = 0
x = -3 or x = -1
-------------------------------
4. A 15 in. TV set also has a width of 12 inches. What is it's height?
15 " is the diagonal
H^2 + 12^2 = 15^2
H^2 = 225 -144 = 81
Height = 9 inches
====================
Cheers,
Stan H.